OpenStudy (anonymous):

When solving this equation by completing the square, which of the following shows the equation with the square completed? x2 - 4x - 32 = 0 A. (x - 4)2 = 36 B. (x - 2)2 = 36 C. (x - 8)2 = 4

OpenStudy (anonymous):

x2 - 4x = 32 x2 - 4x +4 = 32+4 (x-2)2 = 36